MONT 103N – Analyzing Environmental Data Hypothesis Testing Practice Problems April 18, 2012

Directions: For each question, first identify the relevant null and alternative hypotheses. Then carry out the appropriate hypothesis test, estimate the p-value if possible with the information you have, and answer the other questions.

1) The average CO_{2} emissions for a collection of 100 large power plants is \overline{Y} = 2223.1 lb/Mwh, with SD = 211.3 lb/Mwh.

a) If this were a random sample, would there be evidence at the α = .05 level to say that the average coal-fired power plant emits an amount different from 2200 lb/Mwh?

b) This comes from a data set in our textbook given in problem 15 on page 321. Is this a random sample? Does the calculation in part a) really make sense in this setting?

2) In a simple random sample of 2000 Germans in 2001, 840 out of the 2000 thought that all electricity should be generated by “green” energy sources such as wind and solar power. Is there evidence at the α = .05 level to say that at least 40% of all Germans thought the same about sources of electric power?

3) In a study of n = 17 sea star arm lengths, the average was \overline{Y} = 6.8 cm with an SD = .5 cm. Is there evidence at the α = .05 level to say that the population arm length is different from 7 cm? (Note: The short t-table on page 316 of the text has the information you need for this one, but it will not be possible to estimate the p-value with only the information there. Use the t-table demonstrated in class to estimate this.)

4) A medical study compared the resting pulse rates of random, independent samples of 100 smokers and 100 nonsmokers. The smokers had an average pulse rate of \overline{Y} = 86 beats per minute with SD_{1} = 5.4, while the nonsmokers had an average pulse rate of \overline{X} = 80 and SD_{2} = 4.9. Is there evidence at the α = .05 level to say that nonsmokers have a lower average pulse rate?